OpenStudy (anonymous):

complete the square for the equation X^2+2x-3=0

OpenStudy (anonymous):

\[x ^{2} + 2x -3\]

OpenStudy (anonymous):

\[= x ^{2} +3x -x - 3\]

OpenStudy (larseighner):

Don't complete the square and you can factor.

OpenStudy (anonymous):

\[= x(x+3) - 1(x+3)\]

OpenStudy (anonymous):

\[= (x-1)(x+3)\]

OpenStudy (larseighner):

yes

OpenStudy (larseighner):

This is a silly exercise, because completing the square is unnecessary to solve it, since the factors are clearly evident on inspection.

OpenStudy (anonymous):

i dont get it :/

OpenStudy (anonymous):

you don't? where do you need help? tell me

OpenStudy (larseighner):

You can complete the square by adding 4 to both sides That gives you x^2+2x-3 +4 = +4 x^2+2x + 1 = +4

OpenStudy (larseighner):

You should recognize that both sides are perfect squares now, to make it clear: x^2+2x + 1 = +4 (x + 1)(x + 1) = 4 There are several ways to go here. One is to take the root of both sides: \[ \large \sqrt{(x+1)^2} = \pm 2 \] \[ \large x+1 = \pm 2 \] That gives you x+3 and x-1 But this is useless busy work because you could have factored the original expression to begin with.

OpenStudy (larseighner):

Another way: x^2+2x-3 +4 = +4 x^2+2x+1= +4 (x+1)^2 = +4 (x+1)^2 - 4 = 0 This is the difference of two squares, so you should see it factors to [(x+1) + 2][(x+1) - 2] = 0 (x+3)(x-1) = 0

OpenStudy (larseighner):

@Maravega33 Are you still with us? Does one of these explanations make more sense to you?

OpenStudy (anonymous):

Use the formula for CTS if you get lost.